OpenStudy (anonymous):

Find the midpoint of the line segment whose end points are given. (3, 1) and (7, 6)

OpenStudy (cwrw238):

midpoint = [ (x1 + x2) / 2 , (y1 + y2) / 2] where the 2 points are (x1,y1) and (x2,y2)

OpenStudy (anonymous):

y-1=-5/4(x-3)?

OpenStudy (anonymous):

midpoints is -5/4

OpenStudy (cwrw238):

there are 2 questions here

OpenStudy (cwrw238):

the midpoint required in first question = (3+7) / 2 , (1 + 6)/2 = (5, 7/2)

OpenStudy (anonymous):

k so 5,7/2 is the answer i got it right! thank you

OpenStudy (cwrw238):

yw